OpenStudy (anonymous):

2x^4 + 5x^3 + 3x^2 + 15x – 9 = 0 Find all roots and factor completely

OpenStudy (angela210793):

2x^4 + 5x^3 + 3x^2 + 15x – 9 = 0 (x+3)(2x-1)(x^2+3)=0 x=1/2 or x=-3

OpenStudy (anonymous):

can you explain how u factored pls

OpenStudy (anonymous):

we should get 4 roots right...?

OpenStudy (anonymous):

two roots and then factored completely

OpenStudy (anonymous):

angela the wolfram doesnt show steps can u help me out pls i need to know this for a test

OpenStudy (angela210793):

yea Wolfram says thr r two complex solutions too tht come from (x^2+3)=0

OpenStudy (anonymous):

but how do you get (x^2 + 3)= 0

OpenStudy (angela210793):

(x+3)(2x-1)(x^2+3)=0 is zero only whn (x+3)=0---> x=-3 or (2x-1)=0--->x=1/2 or x^2-3=0---> x^2=3--->x=+-sqr3

OpenStudy (anonymous):

o ok thanks

OpenStudy (angela210793):

welcome....(sorry for factorization) :/

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

(2x^4 +3x^2 -9) + (5x^3 + 15x) = 0 (x^2 + 3)(2x^2 + 3) + (5x)(x^2 + 3) = 0 (x^2 + 3)(2x^2 +3 + 5x) = 0 (x^2 +3)(x + 3)(2x - 1) = 0 x + 3 = 0 x = -3 2x - 1 = 0 x = 1/2